OpenStudy (anonymous):

Find the real solutions of the equation 4(x + 1)^2 + 21(x + 1) + 5 = 0

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so this is an equation that is reducible to a quadratic... make a substitution of u = (x + 1) so the equation is \[4u^2 + 21u + 5 = 0\] can you solve this..?

OpenStudy (anonymous):

I'm not sure. Would I just plug it into the quadratic formula?

OpenStudy (campbell_st):

you can here here is a cool method to show your teacher... for a quadratic \[ax^2 + bx + c = 0\] find \[x \times c =\] so you need \[4 \times 5 = 20\] find the factors of 20 that add to b... or 21 in your question...what would you say..?

OpenStudy (anonymous):

20 * 1

OpenStudy (anonymous):

For where you put x * c did you mean a * c?

OpenStudy (campbell_st):

yes... so the factors of 20 are 20 and 1... to they add to given the coefficient of the middle term.. so next is (au + factor 1)(au + factor 2) (4u + 20)(4u + 1) ------------------------- --------------- a 4 now remove any common factors from the binomials 4(u + 5)(4u + 1) --------------- 4 cancel the common factor and you get (u + 5)(4u + 1) = 0 so now you can find the values of u...that make the equation true what are they..?

OpenStudy (anonymous):

u = -5 u = -1/4

OpenStudy (campbell_st):

correct... so now use the substitution again u = x + 1 so x + 1 = -5 and x + 1 = -1/4 hope it helps

OpenStudy (anonymous):

Thanks :)